Recent Log Entries

FSD Masking Tool: Successful System Masking for: BeamLossMonitor

Log Number: 3829783Logbook(s): Tags: Body: System Masking for: BeamLossMonitor
User: mmerz
Areas: All

BeamLossMonitor Devices:
ILM0I07
ILM0L01A
ILM0L01B
ILM0L02
ILM2D00A
ILM2D00B
ILM0L04A
ILM0L04B
ILM0L05
ILM0L06A
ILM0L08
ILM0R01
ILM4D00
ILM0R03
ILM0R04
ILM0R06
ILM0R08
ILM0R09
ILM1L01
ILM1L02
ILM6R08
ILM3S00
ILM7S00
ILM9S00
ILM5S10
ILM5E03
ILM7S01
ILM9A01
ILM3E02
ILM5A01
ILM1A04
ILM1A06
ILM3R10
ILM9R07
ILM1R09
ILM2L03
ILM5R10
ILM5R05
ILM2L01
ILM2T09
ILM8T09
ILM4C01
ILM4C02
ILM4C03
ILM1C01<

Hall A FFB bpms in Normal SEE mode

Log Number: 3829780Logbook(s): Body: 

Taking a look at Hall A SEE bpms on IOCSE9 after reading email from Jay (indicating these bpms were not showing positions when Hall A was receiving .3microamps cw this weekend). While Ops had a Rf trip I looked at IOCSE9 while in SEE normal mode and its live_display did not look correct. Terry Carlino had to try 3 times before he could get channel access open on IOCSE9. Once channel access was open I performed a Check_VMIC and then after that IOCSE9's Live_display looked correct.

LEM South Linac

Log Number: 3829778Logbook(s): Tags: Body: LEM 9.6 was successfully run for the South Linac on Tue Aug 18 07:27:57 2020. The application phase took 0:00:11 to complete.

LEM South Linac - failure

Log Number: 3829777Logbook(s): Tags: Body: LEM 9.6 was unsuccessfully run for the South Linac on Tue Aug 18 07:27:23 2020. The operation terminated after 0:00:07 with the following results.
Cavity summarySuccess: 0 Abort: 200 Fail: 0

FudgeIt

Log Number: 3829776Logbook(s): Tags: Body: 

FudgeIt 4.0.1 changed the South Linac fudge factor from 0.957 to 0.961.

FSD Masking Tool: Successful System Masking for: CryoModule

Log Number: 3829772Logbook(s): Tags: Body: System Masking for: CryoModule
User: mmerz
Areas: All

CryoModule Devices:
0L02
0L04
0L03
1L14
1L03
1L02
1L04
1L07
1L05
1L06
1L08
1L11
1L09
1L10
1L12
1L15
1L13
1L16
1L19
1L17
1L18
1L20
1L21
1L22
1L23
1L24
1L25
1L26
2L14
2L03
2L02
2L04
2L07
2L05
2L06
2L08
2L11
2L09
2L10
2L12
2L15
2L13
2L16
2L19
2L17
2L18
2L20
2L21
2L22
2L23
2L24
2L25
2L26


RF Recovery. No Tunnel Entries.

Log Number: 3829769Logbook(s): Entry Makers: beneschBody: 

To quote the PD's shift plan for the DAY crew:

Deliver beam as requested to all halls until 0745. Shut off beam to halls so they've cooled off a bit when RadCon arrives at 0800 to survey A and C. Leave everything on in CEBAF. No tunnel entries planned, just Halls.

LEM South Linac

Log Number: 3829768Logbook(s): Tags: Body: LEM 9.6 was successfully run for the South Linac on Tue Aug 18 07:14:24 2020. The application phase took 0:02:27 to complete.

2L07-7 filament error

Log Number: 3829767Logbook(s): Entry Makers: friersonBody: 

This prevented the zone from going back into filaments. We decided to bypass and reLEM

CREW CHIEF SIGN IN

Log Number: 3829765Logbook(s): Body: 
THOMAS JEFFERSON NATIONAL ACCELERATOR FACILITY
ACCELERATOR OPERATIONS CREW CHIEF SIGN IN FORM

Date : 08-18-20
Time : 07:11:26
Day : Tuesday
Shift : Day

Crew Chief : Carlino
Program Deputy: Benesch
Team : X
Arm : Gauthier
Designated SSO: Mock
LERF Operator:
Operators : Gauthier, Mock

This stamp certifies that the above Crew Chief has read/reviewed/received the following:

The following cavities have been bypassed: 2L07-7

Log Number: 3829763Logbook(s): Tags: Body: 

The following cavities have been bypassed:
2L07-7
Username: frierson
Date/Time: 08-18-20_07:09:46
RF Status: Bad
Tuner Status: Good
Comment: Bypassed - stepped off +100k - Micro filament enable fault
Special Instructions:

Pages